OpenStudy (anonymous):

A bacteria culture starts with 1000 bacteria and the population triples after 3 hours. Find an expression for the number y(t) of bacteria after t hours. What I have so far is: y(t)=1000e^(?)

OpenStudy (anonymous):

in not familiar with this type of problem, but go to http://answers.yahoo.com/question/index?qid=20101014073636AAw0Trf site. it has an example to ur question

OpenStudy (anonymous):

I was looking at the same problem earlier and I thought I had the answer but it kept saying it was wrong. So I thought I typed in wrong and tried different ways of typing it but it didn't work. So I was hoping I could figure out the answer to see if I was completely wrong or wrote in wrong.

OpenStudy (anonymous):

that problem is after 6 hours, urs is after 3 hours.. did u change the time

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y(t) = 1000*e^((ln(3)/3)*t) did u put it like that?

OpenStudy (anonymous):

Ah thank you! It was a typo on how I put it in.

OpenStudy (wolf1728):

How's this? 0 hours population = 1,000 3 hours population = 3,000 6 hours population = 9,000 So population = 1,000*3^(t/3) 0 hours = 1,000*3^0 = 1,000^1 = 1,000 3 hours = 1,000*3^(t/3) 3 hours = 1,000*3^1 = 1,000 * 3 = 3,000 6 hours = 1,000*3^(t/3) 6 hours = 1,000*3^(6/3) 6 hours = 1,000*3^(2) = 1,000*3² = 9,000

OpenStudy (anonymous):

welcome :)